the area of a rectangle is the product of it's length and width and equals 1062. if the length is 5 more than 3 times the width find the width of the rectangle
Answer by RAY100(1637)

Area = Length * Width = (3w+5)*(w) = 1062
.
3w^2 +5w-1062 = 0
.
use quadratic formula to factor
.
a= 3,,,,b=5,,,,,,c= -1062
.
w= [ -(5) +/- sqrt { (5)^2 -4 (3)(-1062)} ] /2(3)
.
w= [ -5 +/- sqrt{ 12769} ] / 6
.
w=( -5 +/- 113) / 6
.
w= 18, -19 2/3 (not reasonable)
.
checking
.
A=(3*18+5) *(18) = 59*18 = 1062